OpenStudy (anonymous):

How to solve cos10x = 0 and sin10x = 0

OpenStudy (amoodarya):

sin x= sin a x=a+2kpi x=pi-a+2kpi so sin 10x=0 10x=0 +2kpi 10x=pi-0+2kpi so 10x=0=kpi x=kpi/5